OpenStudy (anonymous):

Solve for x: (x - 9)^2 = 1

jimthompson5910 (jim_thompson5910):

(x - 9)^2 = 1 x - 9 = sqrt(1) or x - 9 = -sqrt(1) x - 9 = 1 or x - 9 = -1 Now solve each individual equation for x